The frequency of the second overtone of a closed pipe
is in unison with the third overtone of an open pipe
other ratio of lengths of the pipe is
(a) 10: 8
(b) 5: 8
(c) 7: 4
(d) 4: 7

Grade:Upto college level

2 Answers

Arun
3 years ago

3rd overtonr of closed = 2nd harmonic of open

7\left [ \frac{V}{4l_{c}} \right ]= 2\left [ \frac{V}{l_{0}} \right ]

\frac{l_{c}}{l_{0}}= \frac{7}{4}

 

Closed organ pipe -

A closed organ pipe is a cylindrical tube having an air colume with one end closed.

- wherein

Condition of constructive interference

l=\left ( 2n+1 \right )\frac{\lambda }{4}    

Vikas TU
3 years ago
Third overtone of closed organ pipe means seventh harmonic. Given
(f7)closed=(f4)open
7(ν/4IC)=4(ν/2I0)
lC/l0=7/8